# High-Temperature Fermionic Gibbs States are Mixtures of Gaussian States

### v1

@@ Abstract
Efficient simulation of a quantum system generally relies on structural properties of the quantum state. Motivated by the recent results by Bakshi et al. on the sudden death of entanglement in high-temperature Gibbs states of quantum spin systems, we study the high-temperature Gibbs states of bounded-degree local fermionic Hamiltonians, which include the special case of geometrically local fermionic systems. We prove that at a sufficiently high temperature that is independent of the system size, the Gibbs state is a probabilistic mixture of fermionic Gaussian states. This forms the basis of an efficient classical algorithm to prepare the Gibbs state by sampling from a distribution of fermionic Gaussian states.
@@ 1 Introduction
Main results. We show that at sufficiently high temperatures independent of the system size, the Gibbs state of a geometrically local fermionic Hamiltonian is a classical distribution over a special class of fermionic Gaussian states and can be efficiently sampled. Gaussian states are the ground states and Gibbs states of 2-local fermionic Hamiltonians (free fermions); they are structurally simple, efficiently representable on classical computers, and are often considered the fermionic analogue of product states [18].
@@
[…]
@@
One may also ask whether the high-temperature fermionic Gibbs states can be represented as a mixture of unentangled states, as done in [5]. Here we present an informal argument, which is independent of the rest of the paper, to show that the fermionic superselection rule precludes such a representation. The fermionic superselection rule forbids superposition between even and odd parity sectors. Mathematically, for a system consisting of modes $\gamma_{1},\gamma_{2},\cdots,\gamma_{2n}$, this means any valid quantum state has to commute with the parity operator $i^{n}\gamma_{1}\gamma_{2}\cdots\gamma_{2n}$. If a fermionic system is decomposed into subsystems $S_{1},S_{2},\cdots,S_{\ell}$ with parity operators $P_{1},P_{2},\cdots,P_{\ell}$ respectively, and a quantum state is unentangled across these subsystems, then it needs to simultaneously commute with parity operators $P_{j}$. If a Gibbs state can be expressed as a mixture of such states, then it also needs to commute with all operators $P_{j}$ by linearity. Since the Gibbs state is proportional to $e^{-\beta H}$, we also need $[H,P_{j}]=0$. This requirement is obviously not satisfied by all Hamiltonians, since a single quadratic term across two subsystems $j,k$ already fails to commute with both $P_{j}$ and $P_{k}$.
Technical highlights. We follow the framework developed for spin systems [5]. Here we highlight the necessary modifications to accommodate the fermionic anti-commutation relations, which lead to an adaptive choice of the matching $M^{(i)}$ in the Gaussian states.
@@
We further show that at high temperatures, each $M_{i}$ is a probability distribution over operators of a simple form. At a high level, we use this simple form to randomly sample a simple operator from $M_{1}M_{1}^{\dagger}$, isolate (pin out) the first site $\gamma_{1}$, and move it to the left of $M_{2n}\dots M_{2}M_{2}^{\dagger}...M_{2n}^{\dagger}$. This procedure leaves us with an operator that acts only on the remaining Majorana operators $[2,3...,2n]$, allowing us to recursively apply similar techniques to isolate subsequent operators.
@@
Open problems. Our results implies that high-temperature Gibbs states of geometrically local fermionic Hamiltonians—and more generally, bounded-degree local fermionic Hamiltonians— is a probability distribution over Gaussian states (mixture of Gaussian states) and can be efficiently simulated on classical computers. It remains to be seen whether our techniques can be extended to rule out quantum advantage in other fermionic settings. In particular, [20, 21] showed that quantum computers can efficiently prepare fermionic Gibbs states at any temperature when the non-quadratic interaction is sufficiently weak. An interesting open question is whether these Gibbs states can also be expressed as mixtures of Gaussian states, thereby enabling efficient classical simulation. Alternatively, it may be shown that in this regime the Gibbs state is not a mixture of Gaussian states, in which case the algorithmic framework developed in [5] would no longer apply. Besides, it is also interesting to investigate whether expressing fermionic Gibbs states as mixture of Gaussian states is related to the performance of quantum Monte Carlo methods in the presence of a sign problem [22, 23, 24]. We can see that the results presented in this work open up numerous directions for future investigation.
@@
Overview. In this paper, we will construct two algorithms, a structural algorithm and a sampling algorithm. The two algorithms rely on different, but both constant, temperature bounds. The structural algorithm does not classically prepare the Gibbs state as a probability distribution of Gaussian states, but its correctness implies the structural result that the Gibbs state lies in the convex hull of Gaussian states.
@@
Organization. The rest of the paper is organized as follows: In Section 2 we provide the necessary definitions of Majorana operators and the Hamiltonians that we are going to study. In Section 3 we break the Gibbs state into a product of factors, each of which can be efficiently sampled. In Section 4 we present a sampling algorithm over unnormalized fermionic Gaussian states that produces the Gibbs state in expectation value, which proves the structural result that the Gibbs state is a mixture of Gaussian states. In Section 5 we modify the algorithm in the previous section to obtain an efficient algorithm that outputs random normalized fermionic Gaussian states whose expectation value is the Gibbs state.
@@ 2 Problem Setup
We use $[2n]$ to denote the set $\{1,2,...,2n\}$. For a subset $S\subseteq[2n]$, we use $|S|$ to denote the number of elements in $S$, and $S^{c}$ to denote the complement set $[2n]\backslash S.$ For two subsets $A,B\subseteq[2n]$, we use the notation $A-B=\{a\in A|\,a\notin B\}$, so that the operation is well-defined even when $B\not\subseteq A$.
@@
We use Majorana string to refer to a product of an even number of distinct Majorana operators, that is, $G=\alpha\gamma_{i_{1}}\cdots\gamma_{i_{k}}$, with a complex coefficient $\alpha$ and even $k$. We define its support to be $\text{supp}(G):=\{i_{1},...,i_{k}\}$. We define $\mathcal{M}$ to be the set of all $\mathcal{R}$-local Majorana strings with coefficient $i^{\mathcal{R}/2}$, which is used to ensure that the term is Hermitian, that is, $\left(i^{\mathcal{R}/2}\gamma_{i_{1}}...\gamma_{i_{\mathcal{R}}}\right)^{\dagger}=i^{\mathcal{R}/2}\gamma_{i_{1}}...\gamma_{i_{\mathcal{R}}}$. Moreover, we use $\mathcal{M}^{*}$ to denote the set of all Majorana strings with coefficient in {±$\{\pm 1,\pm i\}$ (so either Hermitian or anti-Hermitian Majorana strings).
@@
where $0\leq\lambda_{a}\leq 1$, $|\text{supp}(G_{a})|$ is even and $\leq\mathcal{R}$, and any local term $G_{a}$ has overlapping support with at most $d$ other terms, i.e. |$|\{b\,|\,\text{supp }(G_{a})\cap\text{supp}(G_{b})\neq\emptyset\}|\leq d$.
@@ 3 Telescoping and Sampling
In the following, we will simplify Eq. (9) by observing that the commutator $[H^{S_{i}},\Gamma_{j,t}]$ is non-zero only on the local terms where $H^{S_{i}}$ and $\Gamma_{j,t}$ overlap. Define $T_{1,j}$ as the set of local terms of $H^{S_{i}}$ whose support overlaps with $\Gamma_{j,t}$:
@@
respectively, with $\Gamma_{j^{\prime},t}$ accordingly set to either $G_{a}\Gamma_{j,t}$ or $-\Gamma_{j,t}G_{a}$. Now, $\mu_{j^{\prime},t+1}$ is the distribution obtained by first sampling $(c_{j,t},\Gamma_{j,t})$ from $\mu$, and then sampling a random value uniformly from $T_{1,j}$ with probability $P=\frac{2td}{2td+(d+1)}$ or from $T_{2}$ with probability $1-P=\frac{d+1}{2td+(d+1)}$. Note that $P\geq\frac{td}{(t+1)d}$ and $1-P\geq\frac{d+1}{2(t+1)d}$. Now, if $(j,t+1)$ corresponds to a value chosen from $T_{1,j}$, we have
@@ 4.1 Overview
[…]
@@
Sampling one of these two terms, setting $\Gamma_{1}$ to be ±$\pm\Gamma_{1}$ and $\sigma_{1}=(I\pm i\gamma_{i}\gamma_{k})\sigma_{0}$ consistently, we have extracted site $i$ and completed one step of the algorithm! Since eventually every site must be pinned, by the end of the algorithm $\Gamma_{2n-1}$ will have trivial support, i.e. will be a scalar multiple of the identity. At this point, $(I+\alpha_{2n-1})\Gamma_{2n-1}$ is a scalar, so defining $\sigma_{2n}\coloneq(I+\alpha_{2n-1})\Gamma_{2n-1}\sigma_{2n-1}$ yields an unnormalized Gaussian state, as desired.
@@ 4.2 Algorithmic Description
      If there are two indices $k,l\in S_{j}^{c}$ such that $k,l\in\text{supp}\left(\Gamma_{j+1}\right)$, i.e. $\Gamma_{j+1}=\Gamma^{-}_{j+1}\cdot(i\gamma_{k}\gamma_{l})$ for some operator $\Gamma^{-}_{j+1}$ whose support is $\text{supp}(\Gamma_{j+1})\backslash\{k,l\}$,:
@@ 4.3.1 Notation
We will also define a filtration $\{\mathcal{F}_{j}^{(x)}\}$. Each $\mathcal{F}_{j}^{(x)}$ is a $\sigma$-algebra that represents all probabilistic events that can be expressed in terms of the randomness in the algorithm until the end of step $(x)$ of the $j$th iteration. In other words, $\mathcal{F}_{j}^{(x)}$ represents the information about the outcomes of random choices up until this point. In particular, we abbreviate $\mathcal{F}_{j}=\mathcal{F}_{j}^{(g)}$, the randomness until iteration $j-1$ ends. We also extend our definition of $\mathcal{F}_{j}$ to $j=0$ by defining $\mathcal{F}_{0}$ to be the trivial $\sigma$-algebra, since at the beginning of the algorithm no stochastic choices have occurred yet. We may then define the conditional expectation $\mathbb{E}\left[X|\mathcal{F}_{j}^{(x)}\right]$ to be the average of $X$ over any uncertainty not contained in $\mathcal{F}_{j}^{(x)}$. It may be thought of as the best guess for $X$, conditioned on the knowledge of whether any event in $\mathcal{F}_{j}^{(x)}$ occurred or did not occur. For instance,

### v2

@@ Abstract
Efficient simulation of a quantum system generally relies on structural properties of the quantum state. Motivated by the recent results by Bakshi et al. on the sudden death of entanglement in high-temperature Gibbs states of quantum spin systems, we study the high-temperature Gibbs states of bounded-degree local fermionic Hamiltonians, which include the special case of geometrically local fermionic systems. We prove that at a sufficiently high temperature that is independent of the system size, the Gibbs state is a probabilistic mixture of fermionic Gaussian states. This forms the basis of an efficient classical algorithm to prepare the Gibbs state by sampling from a distribution of fermionic Gaussian states. As a contrasting example, we show that high-temperature Gibbs states of the Sachdev-Ye-Kitaev (SYK) model are not convex mixtures of Gaussian states.
@@ 1 Introduction
1.1 Main results
We show that at sufficiently high temperatures independent of the system size, the Gibbs state of a geometrically local fermionic Hamiltonian is a classical distribution over a special class of fermionic Gaussian states and can be efficiently sampled. Gaussian states are the ground states and Gibbs states of 2-local fermionic Hamiltonians (free fermions); they are structurally simple, efficiently representable on classical computers, and are often considered the fermionic analogue of product states [38].
@@ 1.1 Main results
It is worth noting that in the above theorem, the partial matching $M^{(i)}$ can vary arbitrarily with $i$. This leads to the key conceptual and technical distinction between the fermionic and spin cases. In the spin setting [1], high-temperature Gibbs states are classical distributions over product states, all defined with respect to a fixed tensor product structure (e.g., qubit $1\otimes$ qubit $2\otimes\cdots\otimes$ qubit $n$). In contrast, in the fermionic case, each Gaussian state $\rho^{(i)}$ corresponds to a different matching $M^{(i)}$ of Majorana operators. While it is possible to reorder the Majorana operators so that a specific $\rho^{(i)}$ becomes a product state under the Jordan–Wigner transformation, there is no obvious way that transforms all $\rho^{(i)}$ into product states simultaneously.
@@
One may also ask whether the high-temperature fermionic Gibbs states can be represented as a mixture of unentangled states, as done in [1]. Here we present an informal argument, which is independent of the rest of the paper, to show that the fermionic superselection rule precludes such a representation. The fermionic superselection rule forbids superposition between even and odd parity sectors. Mathematically, for a system consisting of modes $\gamma_{1},\gamma_{2},\cdots,\gamma_{2n}$, this means any valid quantum state has to commute with the parity operator $i^{n}\gamma_{1}\gamma_{2}\cdots\gamma_{2n}$. If a fermionic system is decomposed into subsystems $S_{1},S_{2},\cdots,S_{\ell}$ with parity operators $P_{1},P_{2},\cdots,P_{\ell}$ respectively, and a quantum state is unentangled across these subsystems, then it needs to simultaneously commute with parity operators $P_{j}$. If a Gibbs state can be expressed as a mixture of such states, then it also needs to commute with all operators $P_{j}$ by linearity. Since the Gibbs state is proportional to $e^{-\beta H}$, we also need $H,P_{j}]=0$. This requirement is obviously not satisfied by all Hamiltonians, since a single quadratic term across two subsystems $j,k$ already fails to commute with both $P_{j}$ and $P_{k}$.
In addition to our structural and sampling results for bounded-degree local fermionic Hamiltonians, we present a complementary negative result for the Sachdev-Ye-Kitaev (SYK) model. As a contrasting example, we find that high-temperature Gibbs states of the Sachdev-Ye-Kitaev (SYK) model are not mixtures of Gaussian states, as stated in Theorem 1.3. Note that while each interaction term in SYK is $q$-local, the model has unbounded degree thus Theorem 1.1 and Theorem 1.2 do not apply. This shows that the structural phenomena we establish for bounded-degree local fermionic Hamiltonians break down in strongly interacting systems, such as SYK. The proof builds on prior work [15, 17, 13] that studies the ground state of the SYK model.
Theorem 1.3 (Informal version of Theorem 6.5)
Let $q\geq 12$ be an even integer, and let $SYK_{q}$ denote the distribution over $q$-local Hamiltonians with unbounded degree on $2n$ Majorana modes defined by
$H=i^{q/2}\sum_{1\leq k_{1}<...<k_{q}\leq 2n}J_{k_{1}\dots k_{q}}\gamma_{k_{1}}\dots\gamma_{k_{q}}.$ (3)
where each coefficient $J_{k_{1}\dots k_{q}}$ is an independent standard Gaussian variable with mean 0 and variance $1$. For $\epsilon>0$ and any inverse temperature $\beta=\Omega(n^{-3q/4+1+\epsilon})$, with high probability over $H\sim\text{SYK}_{q}$ there exists a $0<\beta^{\prime}<\beta$ for which the Gibbs state
$\rho_{\beta^{\prime}}=\frac{e^{-\beta^{\prime}H}}{\text{tr}\left(e^{-\beta^{\prime}H}\right)}$ (4)
is not a mixture of Gaussian states.
Note that the SYK model in physics and computer science literature is typically normalized by a coefficient either scaling as $n^{(q-1)/2}$ or $n^{q/2}$. With either choice of normalization, the above theorem implies as a corollary that for any constant temperature $\beta^{-1}$, there exists a $0<\beta^{\prime}<\beta$ for which $\rho_{\beta^{\prime}}$ is not a mixture of Gaussian states.
1.2 Technical highlights.
We follow the framework developed for spin systems [1]. Here we highlight the necessary modifications to accommodate the fermionic anti-commutation relations, which lead to an adaptive choice of the matching $M^{(i)}$ in the Gaussian states.
@@
We further show that at high temperatures, each $M_{i}$ is a probability distribution over operators of a simple form. At a high level, we use this simple form to randomly sample a simple operator from $M_{1}M_{1}^{\dagger}$, isolate (pin out) the first site $\gamma_{1}$, and move it to the left of $M_{2n}\dots M_{2}M_{2}^{\dagger}...M_{2n}^{\dagger}$. This procedure leaves us with an operator that acts only on the remaining Majorana operators [2,3,2n][2,3...,2n], allowing us to recursively apply similar techniques to isolate subsequent operators.
@@
To complement our structural results for bounded-degree local Hamiltonians, we also establish that the Gibbs states of the Sachdev–Ye–Kitaev (SYK) model are not mixtures of Gaussian states at sufficiently high temperatures. Our argument is based on comparing energies achieved by the Gibbs states and Gaussian states. In particular, we show that for an SYK Hamiltonian $H$, the Gibbs state $\rho_{\beta}$ achieves a better energy than any Gaussian state can. We achieve this by calculating the derivative of $\text{tr}(\rho_{\beta}H)$ with respect to $\beta$ at $\beta=0$ and upper bounding the corresponding second derivative. Combining this approach with known upper bounds for the energy of Gaussian states, we conclude that for $q\geq 12$, the SYK Gibbs state cannot be convex-Gaussian.
1.3 Contribution and open problems
In this work we prove that the high-temperature fermionic Gibbs state is a mixture of fermionic Gaussian states. By modifying the sampling procedure used in this proof, we also provide an efficient algorithm to sample normalized Gaussian state density matrices whose expectation value yields the target Gibbs state. Our result rigorously demonstrates the absence of the fermionic sign problem at high temperature, and sharpens the boundary between quantum advantage and classical simulability for fermionic systems. As a contrasting example, we find that high-temperature Gibbs states of the Sachdev–Ye–Kitaev (SYK) model are not mixtures of Gaussian states, showing that the structural phenomena for bounded-degree fermionic Hamiltonians break down in strongly interacting systems.
To understand the limitations of simulating quantum systems on classical computers, one needs to look beyond exactly solvable models, and explore examples that are at the boundary between hard and easy systems. These examples include weakly interacting spin systems [5], the quantum impurity model [7], and quantum systems at high temperature. Recently, significant progress has been made in the rigorous analysis of the high-temperature regime, leading to provable algorithms for sampling from the Gibbs state [45, 1] and for computing the partition function in the presence of long-range interactions [34]. Our result represents a step towards better understanding of fermionic systems at high temperature and our techniques provide insights for generalizing results for spin systems to the fermionic setting.
@@ 1.3 Contribution and open problems
We can see that the results presented in this work open up numerous directions for future investigation. Our results implies that high-temperature Gibbs states of geometrically local fermionic Hamiltonians—and more generally, bounded-degree local fermionic Hamiltonians—is a probability distribution over Gaussian states and can be efficiently simulated on classical computers. It remains to be seen whether our techniques can be extended to rule out quantum advantage in other fermionic settings. In particular, [42, 36] showed that quantum computers can efficiently prepare fermionic Gibbs states at any temperature when the non-quadratic interaction is sufficiently weak. An interesting open question is whether these Gibbs states can also be expressed as mixtures of Gaussian states, thereby enabling efficient classical simulation. Alternatively, it may be shown that in this regime the Gibbs state is not a mixture of Gaussian states , in which case the algorithmic framework developed in [1] would no longer apply, suggesting the possibility of quantum advantages.
It is also natural to consider how the techniques in this work and in [1] relate to quantum Monte Carlo methods such as determinant quantum Monte Carlo and auxiliary-field quantum Monte Carlo [3, 46, 47, 23]. A key difference in the algorithm we use is that we only sample positive semidefinite density matrices whose expectation value yields the Gibbs state, whereas quantum Monte Carlo methods typically sample wavefunctions, e.g., in the form of Slater determinants, together with weights that may have complex phases. This key difference is one reason why our algorithm does not suffer from the fermionic sign problem [41, 19, 35]. It is of interest to investigate whether the structural result of expressing fermionic Gibbs states as mixture of Gaussian states is useful for ameliorating the sign problem. The connection to Monte Carlo methods for stoquastic Hamiltonians [39, 6, 8, 11, 40], which is typically free from the sign problem, is another direction to be explored.
Additionally, it is interesting to explore whether our results could inspire new learning algorithms for fermionic systems. In classical machine learning, mixtures of Gaussian distributions are often learned using the Expectation–Maximization (EM) algorithm [2, 12] —an iterative method which monotonously increases the likelihood and differs significantly from existing approaches for learning fermionic states [27, 26, 25]. Given our structural result that high-temperature fermionic Gibbs states can be expressed as mixtures of fermionic Gaussian states, it is natural to ask whether the EM algorithm can be adapted to efficiently learn such states.
1.4 Overview and organization
@@
In this paper, we will construct two algorithms, a structural algorithm and a sampling algorithm. The two algorithms rely on different, but both constant, temperature bounds. The structural algorithm does not classically prepare the Gibbs state as a probability distribution of Gaussian states, but its correctness implies the structural result that the Gibbs state lies in the convex hull of Gaussian states.
@@
Organization. The rest of the paper is organized as follows: In Section 2 we provide the necessary definitions of Majorana operators and the Hamiltonians that we are going to study. In Section 3 we break the Gibbs state into a product of factors, each of which can be efficiently sampled. In Section 4 we present a sampling algorithm over unnormalized fermionic Gaussian states that produces the Gibbs state in expectation value, which proves the structural result that the Gibbs state is a mixture of Gaussian states. In Section 5 we modify the algorithm in the previous section to obtain an efficient algorithm that outputs random normalized fermionic Gaussian states whose expectation value is the Gibbs state. Lastly, in Section 6 we present the complementary result that high-temperature Gibbs state of the dense SYK model is not a mixture of Gaussian states.
@@ 2 Problem Setup
We use $[2n]$ to denote the set $\{1,2,...,2n\}$. For a subset $S\subseteq[2n]$, we use $|S|$ to denote the number of elements in $S$, and $S^{c}$ to denote the complement set $\S.[2n]\backslash$ S. For two subsets $A,B\subseteq[2n]$, we use the notation $A-B=\{a\in A|\,a\notin B\}$, so that the operation is well-defined even when $B\not\subseteq A$.
@@
We use Majorana string to refer to a product of an even number of distinct Majorana operators, that is, $G=\alpha\gamma_{i_{1}}\cdots\gamma_{i_{k}}$, with a complex coefficient $\alpha$ and even $k$. We define its support to be $\text{supp}(G):=\{i_{1},...,i_{k}\}$. We define $\mathcal{M}$ to be the set of all $\mathcal{R}$-local Majorana strings with coefficient $i^{\mathcal{R}/2}$, which is used to ensure that the term is Hermitian, that is, $\left(i^{\mathcal{R}/2}\gamma_{i_{1}}...\gamma_{i_{\mathcal{R}}}\right)^{\dagger}=i^{\mathcal{R}/2}\gamma_{i_{1}}...\gamma_{i_{\mathcal{R}}}$. Moreover, we use $\mathcal{M}^{*}$ to denote the set of all Majorana strings with coefficient in $\{\pm 1,\pm i\}$ (so either Hermitian or anti-Hermitian Majorana strings).
@@
where $0\leq\lambda_{a}\leq 1$, $|\text{supp}(G_{a})|$ is even and $\leq\mathcal{R}$, and any local term $G_{a}$ has overlapping support with at most $d$ other terms, i.e. $|\{b\,|\,\text{supp }(G_{a})\cap\text{supp}(G_{b})\neq\emptyset\}|\leq d$.
@@ 3 Telescoping and Sampling
In the following, we will simplify Eq. (11) by observing that the commutator $H^{S_{i}},\Gamma_{j,t}]$ is non-zero only on the local terms where $H^{S_{i}}$ and $\Gamma_{j,t}$ overlap. Define $T_{1,j}$ as the set of local terms of $H^{S_{i}}$ whose support overlaps with $\Gamma_{j,t}$:
@@
respectively, with $\Gamma_{j^{\prime},t}$ accordingly set to either $G_{a}\Gamma_{j,t}$ or $\Gamma_{j,t}G_{a}$. Now, $\mu_{j^{\prime},t+1}$ is the distribution obtained by first sampling $(c_{j,t},\Gamma_{j,t})$ from $\mu$, and then sampling a random value uniformly from $T_{1,j}$ with probability $P=\frac{2td}{2td+(d+1)}$ or from $T_{2}$ with probability $1-P=\frac{d+1}{2td+(d+1)}$. Note that $P\geq\frac{td}{(t+1)d}$ and $1-P\geq\frac{d+1}{2(t+1)d}$. Now, if $(j,t+1)$ corresponds to a value chosen from $T_{1,j}$, we have
@@ 4.1 Overview
where $c_{1},\dots,c_{7}\geq 0$ are any values such that $\sum_{i}c_{i}=1$. This expansion may be considered a “sampling” of terms from Equation 16 according to a certain probability distribution. Notice that there is a choice of a probability distribution here, that is specified carefully in the algorithm. Sampling one of the above 7 terms according to the distribution $c_{i}$, we obtain a new expression of the form $(I+\alpha_{1}\Gamma_{1})\sigma_{0}$, where again $\mathbb{E}[M_{1}^{-1}(I+\alpha_{1}\Gamma_{1})\sigma_{0}(M_{1}^{\dagger})^{-1}]=I$. Now, since $\Gamma_{1}\in\mathcal{M}^{*}$, it is Hermitian of anti-Hermitian. However, it can be seen that setting $\Gamma_{1}\coloneq 0$ if it is anti-Hermitian does not change equation 15, since all the anti-Hermitian terms cancel each other out.
@@
Sampling one of these two terms, setting $\Gamma_{1}$ to be $\pm\Gamma_{1}$ and $\sigma_{1}=(I\pm i\gamma_{i}\gamma_{k})\sigma_{0}$ consistently, we have extracted site $i$ and completed one step of the algorithm! Since eventually every site must be pinned, by the end of the algorithm $\Gamma_{2n-1}$ will have trivial support, i.e. will be a scalar multiple of the identity. At this point, $(I+\alpha_{2n-1})\Gamma_{2n-1}$ is a scalar, so defining $\sigma_{2n}\coloneq(I+\alpha_{2n-1})\Gamma_{2n-1}\sigma_{2n-1}$ yields an unnormalized Gaussian state, as desired.
@@ 4.2 Algorithmic Description
[…]
@@ 4.3.1 Notation
We will also define a filtration $\{\mathcal{F}_{j}^{(x)}\}$. Each $\mathcal{F}_{j}^{(x)}$ is a $\sigma$-algebra that represents all probabilistic events that can be expressed in terms of the randomness in the algorithm until the end of step $(x)$ of the $j$th iteration. In other words, $\mathcal{F}_{j}^{(x)}$ represents the information about the outcomes of random choices up until this point. In particular, we abbreviate $\mathcal{F}_{j}=\mathcal{F}_{j}^{(g)}$, the randomness until iteration j1j-1 ends. We also extend our definition of $\mathcal{F}_{j}$ to $j=0$ by defining $\mathcal{F}_{0}$ to be the trivial $\sigma$-algebra, since at the beginning of the algorithm no stochastic choices have occurred yet. We may then define the conditional expectation $\mathbb{E}\left[X|\mathcal{F}_{j}^{(x)}\right]$ to be the average of $X$ over any uncertainty not contained in $\mathcal{F}_{j}^{(x)}$. It may be thought of as the best guess for $X$, conditioned on the knowledge of whether any event in $\mathcal{F}_{j}^{(x)}$ occurred or did not occur. For instance,